OpenStudy (anonymous):

Solve. open parentheses square root of 6 close parentheses to the 8 x power = 216x–

OpenStudy (anonymous):

draw it

OpenStudy (anonymous):

= 21^6x–3

OpenStudy (anonymous):

8x * log(SQRT(6)) = (6x-3) * log(21) Use calculator to find the logs, then re-arrange and solve for x.

OpenStudy (anonymous):

okkk

OpenStudy (anonymous):

?? x = 4

OpenStudy (anonymous):

8x*0.9 = (6x-3)*3.05 7.2x = 18x - 9 => 10.8x = 9 => x = 9/10.8 => x = 0.833

OpenStudy (anonymous):

ohhhk